Given the following information about a product, at Phyllis Simons firm, what is the appropriate setup time? Setup labor cost $50.00 per hour Annual holding cost $13 per unit Daily production1,200 units/day Annual demand 43,500 (290 days each x daily demand of 150 units) Desired lot size 150 units Setup timeminutes (round your response to two decimal places)
Answer #1

EOQ = \sqrt{}(2 * D * S)/ (H ( 1 - d/p))

D = Annual Demand = 43,500 units

S = setup cost = $ 50

H = Annual holding cost = $ 13

d = daily demand = 150 units

p = daily production = 1,200 units

Desired lot size = 150

Setup cost = ((150)2 * 13 * (1-150/1200)) / (2 * 43,500)

= 255,937.5 / 87000 = $ 2.94

Set up time = 2.94 / hourly labour rate = 2.94/50 = 0.0588 hours = 3.53 minutes